OpenStudy (anonymous):

1and 2 are supplementary angles. m 1=x-39, and m 2 = x-9. find the measure of each angle.

OpenStudy (anonymous):

supplementary means their sum is 180 deg so (x-39) + (x-9) = 180 deg x + x - 39 -9 = 180 2x = 180 + 48 x = 228/2 x = 114 so m1 = x-39 = 114-39 = 75 m2 = x - 9 = 114-9 = 105

OpenStudy (anonymous):

wait wrong one hold on